Derive a reductions formula for integral emx/xn dx , where m and n are constants.


1
Expert's answer
2021-05-31T15:18:35-0400

"I=\int \frac{e^{mx}}{x^n}dx\newline
Put \space z=x^{1-n} \implies dz=\frac{1-n}{x^n}dx\newline
I=-\frac{1}{1-n}\int e^{mz^{\frac{1}{1-n}}} dz \newline
Put \space w=m^{1-n}z \implies dw=m^{1-n} \newline
=\frac{m^{n-1}}{1-n}\int e^{w\frac{1}{1-n}} dw \newline
\text{By using gamma function,}
\newline
=-\frac{m^{n-1}}{1-n} (1-n)(-1)^{n-1}\Gamma(1-n, -w^\frac{1}{1-n})
\newline
=m^{n-1}(-1)^{n}\Gamma(1-n, -mx)+C"